OpenStudy (anonymous):

What is -21 1/2 divide by 6 ? A. -2 1/6 Multiplying and Dividing Mixed Numbers B. - 5/12 Multiplying and Dividing Mixed Numbers C. 5/12 Multiplying and Dividing Mixed Numbers D. 2 1/6 Multiplying and Dividing Mixed Numbers

OpenStudy (♪chibiterasu):

Let's see who gets me medal @_@

OpenStudy (charlotte123):

-21 1/2 = -43/2 6 = 6/1 To divide fractions, we turn the second fraction upside-down ^_^ so 6/1 becomes 1/6 :) then we multiply :D -43/2 * 1/6 ---> -43/12 simplified, -43/12 = -3 7/12 ^_^

OpenStudy (anonymous):

maybe charlotte might get it :)

OpenStudy (charlotte123):

oh wait, I think I did a mistake X_X

OpenStudy (anonymous):

i'm phykick

OpenStudy (charlotte123):

study, r u sure those are the right choices? :)

OpenStudy (♪chibiterasu):

http://prntscr.com/223oo9 >_>

OpenStudy (charlotte123):

thats a lot of b's chibi XD

OpenStudy (♪chibiterasu):

user buhbuhbuhbuhbuhbuhbuhbuhbuhbuhbuhbuh o-o

OpenStudy (charlotte123):

http://www.wolframalpha.com/input/?i=-21+1%2F2+divide+by+6 >.> looks like I was right, I just don't see the choice ;-;

OpenStudy (anonymous):

Char you got the right answer but it isn't one of the choices I'm so confused

OpenStudy (anonymous):

yes

OpenStudy (♪chibiterasu):

study be usin mind tricks perpare da torches @_@

OpenStudy (charlotte123):

exactly >.> study, u sure those r the correct choices?

OpenStudy (♪chibiterasu):

CHAR GETS ME MEDAL HUE HUE HUR

OpenStudy (charlotte123):

and if u r, r u sure that the fraction is the right one? XD

OpenStudy (charlotte123):

oh u guys <3 I loe u ;-; ┌(-_-) ┐└(-_-)┘┌(-_-) ┐└(-_-)┘ ♪┏(・o・)┛♪┗ ( ・o・) ┓♪┏ ( ) ┛♪┗ (・o・ ) ┓♪┏(・o・)┛

OpenStudy (anonymous):